OpenStudy (anonymous):

There are 28.35 grams in an ounce and 2.21 pounds in a kilogram. Miriam converted 7 kilograms to ounces, but her answer is not correct. What is the best explanation of her error? mc026-1.jpg Miriam set up the problem correctly but made a multiplication error. Miriam multiplied correctly, but her second fraction should be mc026-2.jpg. Miriam multiplied correctly, but her third fraction should be mc026-3.jpg. Miriam is missing a fraction that is needed to convert the 7 kilograms to pounds. if it says some jpg then copy and paste it in your tool bar will fan and medal

OpenStudy (danjs):

\[\frac{ 7~kg }{ 1 }*\frac{ 1000~g }{ 1~kg }*\frac{ 1~ounce }{ 28.35~g } = \frac{ 7*1000 }{ 28.35 }~ounce\]

OpenStudy (danjs):

Start with given, 7 kg, multiply by ratios of 1 so that units cancel in your favor, left with ounces...

OpenStudy (anonymous):

hmmm

OpenStudy (loser66):

\[\frac{ 7~\cancel {kg }}{ 1 }*\frac{ 1000~\bcancel{g} }{ 1~\cancel{kg} }*\frac{ 1~ounce }{ 28.35~\bcancel {g} } = \frac{ 7*1000 }{ 28.35 }~ounce\]

OpenStudy (anonymous):

idk

OpenStudy (loser66):

neither I :)

OpenStudy (anonymous):

really